OpenStudy (anonymous):

2x+y=-11 3x-4y=11 Please i need the answer and help with understanding how to solve this? much appreciated. the answer choice: 3,5 ***-3,-5 ***-5,-3 ***5,3

OpenStudy (anonymous):

Solve for y. y=-11-2x Substitute into second equation 3x-4(-11-2x)=11 Collect like terms: 11x+44=11 x=-3 substitute into first equation: y=-11-2(-3) y=-5

OpenStudy (anonymous):

2x+y=-11 y=-2x-11 then plug this into the y in the other equation 3x-4(-2x-11)=11 3x+8x+44=11 11x=-33 x=-3 now plug this into either one of the original equations 2(-3) +y=-11 -6+y=-11 y=-5 (-3,-5)

OpenStudy (anonymous):

You can just eliminate the y value. (2x+y=-11) * 4 -> 8x+4y=-44 +3x-4y=11 After adding you get: 11x = -33 so x = -3 Then just plug and chug for y. 2(-3) +y = -11 -6+y=-11 y=-5 So, you get (-3,-5)

OpenStudy (anonymous):

sleek-feathered ones

OpenStudy (anonymous):

lol my calc teacher says plug and chug too

OpenStudy (anonymous):

Thank you so much!!